A boys' club decides to build a cabin. The job can be done by 3 skilled workmen in 4 days or by 5 of the boys in 6 days. How long will it take if all work together? Thanks

Answer by stanbon(57290)

A boys' club decides to build a cabin. The job can be done by 3 skilled workmen in 4 days or by 5 of the boys in 6 days. How long will it take if all work together?
-------------------
Men rates: 3/4 job per day for 3 men = 1/4 job per day for each man
----
Boy rates: 5/6 job per day for 5 men = 1/6 job per day for each boy
-----
Together rate = 1/x job/day
Rate + rate = together rate for 1 man and 1 boy:
(3/4) + (5/6) = 1/x
18x + 20x = 24
38x = 24
x = 12/19 days/job (time to do the job together)
===============
Cheers,
Stan H.